OpenStudy (anonymous):

x + 5y = -2 2x + y = 5 The point of intersection of the lines has a y-coordinate of ___.

OpenStudy (anonymous):

set the equations equal to eachother.

OpenStudy (anonymous):

so i would set it up as x+5y+2=2x+y-5 and then do like terms.

OpenStudy (anonymous):

3x+6y-3

OpenStudy (anonymous):

or wait, im sorry use the elimination or substitution methods.

terenzreignz (terenzreignz):

I have never heard of this approach before. Don't you have to solve the system?

OpenStudy (anonymous):

if you use elimination it would probably be easier.

OpenStudy (anonymous):

x - y = 4 x + y = 8 The x-coordinate of the solution to the system shown is ___. @Hope_nicole elimination

OpenStudy (anonymous):

well, the x coordinate you would find by doing this:

OpenStudy (anonymous):



terenzreignz (terenzreignz):

It would have been easier to just add them outright and get rid of the y instead...

OpenStudy (anonymous):

Find the value of \(x\) in the first equation: \[x=4+y\] Plug that into the second: \[(4+y)+y=8\] \[2y=4\] \[y=2\]

OpenStudy (anonymous):

im looking for the x-coordinate

terenzreignz (terenzreignz):

That's right :) 

OpenStudy (anonymous):

2x + y = -2 x + y = 5 The x-coordinate of the solution to the system shown is ___.

OpenStudy (anonymous):

@brittnicoleee_ , scroll up to the picture i put in, see if it makes sense

OpenStudy (anonymous):

Subtract the second from the first like I showed you here: \(2x+y=-2\) \(-x-y=-5\) ---------- \(x+0=-7\)

OpenStudy (anonymous):

2x + 3y = 6 5x + 2y = 4 Which of the following equations could be the result of multiplication and addition to eliminate a variable in the system of equations?